OpenStudy (anonymous):

Look at the table of values below. a b -1 3 0 5 1 7 2 9 Which equation is represented by the table?

OpenStudy (anonymous):

A. b = a + 4 B. b = a + 5 C. b = a + 7 D. b = 2a + 5

OpenStudy (anonymous):

MEDAL ANYONE ??????????(=

OpenStudy (lukecrayonz):

D :)